# Quantifying Data Contamination in Psychometric Evaluations of LLMs

### v1

@@ Appendix C Additional Results and Analyses
gpt-4o-mini 0.77 0.34 0.93 2.00 0.56
gpt-4o 1.61 0.31 0.97 0.20 1.08
gpt-4.1-nano 0.80 0.34 0.90 0.20 1.58
gpt-4.1-mini 0.78 0.38 0.97 0.30 0.78
gpt-4.1 0.70 0.41 1.00 0.10 0.46
gpt-5-nano 2.04 0.38 0.97 0.60 0.40
gpt-5-mini 0.89 0.34 1.00 0.20 0.44
gpt-5 1.57 0.88 1.00 0.00 0.42
@@
qwen3-14b 1.19 0.22 0.97 0.60 0.51
qwen3-32b 0.86 0.28 1.00 0.90 0.81
qwen3-235b-a22b 0.89 0.34 0.97 0.70 0.38
@@
gemini-2.0-flash 0.83 0.38 0.84 0.50 0.60
gemini-2.5-flash-lite 2.48 0.38 0.89 0.00 1.07
gemini-2.5-flash 0.81 0.31 0.93 0.48 0.67
@@
claude-3.5-sonnet 4.36 0.53 0.97 0.20 0.53
claude-4.5-sonnet 9.94 0.44 1.00 0.67 0.43
@@
llama-3.1-70b 4.00 0.34 0.89 1.52 0.81
llama-3.1-405b 9.23 0.41 0.93 2.18 1.24
Table 4: Psychometric contamination results (MFQ).
Model
Verbatim
Memorization
(AED $\downarrow$)
Key Information
Memorization
(Success Rate $\uparrow$)
Item-Dimension
Mapping
(F1 $\uparrow$)
Option-Score
Mapping
(MAE $\downarrow$)
Target Score
Matching
(MAE $\downarrow$)
@@
gpt-4o-mini 1.02 0.55 0.98 0.23 0.38
gpt-4o 0.75 0.62 0.95 0.07 0.75
gpt-4.1-nano 0.71 0.45 0.89 0.73 1.10
gpt-4.1-mini 0.67 0.53 0.91 0.07 0.35
gpt-4.1 0.64 0.62 0.97 0.00 0.58
gpt-5-nano 1.43 0.45 0.92 0.00 0.08
gpt-5-mini 0.70 0.62 0.95 0.00 0.04
gpt-5 0.71 0.85 1.00 0.00 0.02
@@
qwen3-14b 0.82 0.40 0.87 0.00 0.00
qwen3-32b 1.30 0.53 0.84 0.00 0.00
qwen3-235b-a22b 0.68 0.50 0.88 0.00 0.04
@@
gemini-2.0-flash 0.65 0.68 0.84 0.00 0.46
gemini-2.5-flash-lite 0.75 0.53 0.95 0.00 0.53
gemini-2.5-flash 0.62 0.70 0.95 0.07 0.65
@@
claude-3.5-sonnet 2.19 0.85 1.00 0.00 0.33
claude-4.5-sonnet 0.68 0.80 0.94 0.00 0.33
@@
llama-3.1-70b 1.20 0.62 0.88 0.09 0.41
llama-3.1-405b 6.47 0.62 0.93 1.31 0.92
Table 5: Psychometric contamination results (PVQ-40).
Model
Verbatim
Memorization
(AED $\downarrow$)
Key Information
Memorization
(Success Rate $\uparrow$)
Item-Dimension
Mapping
(F1 $\uparrow$)
Option-Score
Mapping
(MAE $\downarrow$)
Target Score
Matching
(MAE $\downarrow$)
@@
gpt-4o-mini 0.76 0.25 0.92 0.40 0.65
gpt-4o 1.12 0.33 1.00 0.10 0.69
gpt-4.1-nano 0.92 0.21 0.96 0.48 1.56
gpt-4.1-mini 0.80 0.25 0.90 0.20 0.71
gpt-4.1 0.80 0.33 1.00 0.25 0.44
gpt-5-nano 2.98 0.29 0.96 0.00 0.00
gpt-5-mini 1.19 0.46 0.96 0.00 0.00
gpt-5 2.27 0.62 1.00 0.00 0.01
@@
qwen3-14b 2.92 0.29 0.95 0.10 0.06
qwen3-32b 3.35 0.17 0.96 0.00 0.00
qwen3-235b-a22b 0.85 0.25 0.96 0.10 0.06
@@
gemini-2.0-flash 0.93 0.25 0.80 0.20 0.49
gemini-2.5-flash-lite 3.35 0.21 0.96 0.40 0.78
gemini-2.5-flash 0.78 0.29 0.98 0.00 0.42
@@
claude-3.5-sonnet 4.49 0.50 1.00 0.00 0.10
claude-4.5-sonnet 13.51 0.67 0.95 0.00 0.58
@@
llama-3.1-70b 4.14 0.25 0.82 1.25 0.53
llama-3.1-405b 9.22 0.29 1.00 1.88 0.40
Table 6: Psychometric contamination results (SD-3).

### v2

@@ Appendix C Additional Results and Analyses
gpt-4o-mini 0.11 $\pm$ 0.01 0.36 $\pm$ 0.04 0.93 $\pm$ 0.00 2.17 $\pm$ 0.52 0.57 $\pm$ 0.03
gpt-4o 0.37 $\pm$ 0.00 0.32 $\pm$ 0.04 0.97 $\pm$ 0.00 0.17 $\pm$ 0.14 1.12 $\pm$ 0.19
gpt-4.1-nano 0.10 $\pm$ 0.01 0.28 $\pm$ 0.08 0.90 $\pm$ 0.08 0.32 $\pm$ 0.10 1.54 $\pm$ 0.10
gpt-4.1-mini 0.10 $\pm$ 0.01 0.30 $\pm$ 0.09 0.95 $\pm$ 0.05 0.47 $\pm$ 0.38 0.90 $\pm$ 0.09
gpt-4.1 0.32 $\pm$ 0.05 0.41 $\pm$ 0.00 1.00 $\pm$ 0.00 0.10 $\pm$ 0.00 0.43 $\pm$ 0.07
gpt-5-nano 0.12 $\pm$ 0.03 0.32 $\pm$ 0.09 0.99 $\pm$ 0.05 0.47 $\pm$ 0.14 0.31 $\pm$ 0.12
gpt-5-mini 0.22 $\pm$ 0.04 0.44 $\pm$ 0.00 1.00 $\pm$ 0.00 0.20 $\pm$ 0.00 0.43 $\pm$ 0.07
gpt-5 0.31 $\pm$ 0.07 0.81 $\pm$ 0.21 1.00 $\pm$ 0.00 0.03 $\pm$ 0.14 0.39 $\pm$ 0.04
@@
qwen3-14b 0.11 $\pm$ 0.01 0.25 $\pm$ 0.08 0.98 $\pm$ 0.05 0.77 $\pm$ 0.29 0.39 $\pm$ 0.15
qwen3-32b 0.10 $\pm$ 0.01 0.24 $\pm$ 0.09 0.98 $\pm$ 0.05 0.73 $\pm$ 0.38 0.76 $\pm$ 0.22
qwen3-235b-a22b 0.26 $\pm$ 0.03 0.14 $\pm$ 0.04 0.83 $\pm$ 0.19 1.41 $\pm$ 0.99 0.47 $\pm$ 0.19
@@
gemini-2.0-flash-001 0.13 $\pm$ 0.02 0.39 $\pm$ 0.12 0.84 $\pm$ 0.00 0.48 $\pm$ 0.05 0.54 $\pm$ 0.08
gemini-2.5-flash-lite 0.10 $\pm$ 0.02 0.35 $\pm$ 0.12 0.87 $\pm$ 0.06 0.00 $\pm$ 0.00 0.91 $\pm$ 0.02
gemini-2.5-flash 0.26 $\pm$ 0.03 0.30 $\pm$ 0.04 0.94 $\pm$ 0.05 0.49 $\pm$ 0.82 0.79 $\pm$ 0.14
@@
claude-3.5-sonnet 0.15 $\pm$ 0.66 0.52 $\pm$ 0.09 0.97 $\pm$ 0.00 0.21 $\pm$ 0.13 0.64 $\pm$ 0.03
claude-sonnet-4.5 0.95 $\pm$ 0.00 0.46 $\pm$ 0.04 1.00 $\pm$ 0.00 0.66 $\pm$ 0.05 0.41 $\pm$ 0.03
@@
llama-3.1-70b-instruct 0.09 $\pm$ 0.03 0.30 $\pm$ 0.16 0.92 $\pm$ 0.06 1.54 $\pm$ 0.76 0.76 $\pm$ 0.10
llama-3.1-405b-instruct 0.05 $\pm$ 0.19 0.35 $\pm$ 0.16 0.94 $\pm$ 0.03 2.39 $\pm$ 0.31 0.93 $\pm$ 0.31
Average 0.17 0.35 0.95 0.70 0.68
Table 4: Psychometric contamination results for MFQ with 95% confidence intervals from three experimental repetitions.
Semantic Memorization (Similarity $\uparrow$) Key Information Memorization (Success Rate $\uparrow$) Item–Dimension Mapping (F1 $\uparrow$) Option–Score Mapping (MAE $\downarrow$) (MAE $\downarrow$)
@@
gpt-4o-mini 0.21 $\pm$ 0.01 0.54 $\pm$ 0.04 0.97 $\pm$ 0.04 0.25 $\pm$ 0.22 0.40 $\pm$ 0.03
gpt-4o 0.42 $\pm$ 0.03 0.58 $\pm$ 0.04 0.96 $\pm$ 0.04 0.07 $\pm$ 0.00 0.86 $\pm$ 0.06
gpt-4.1-nano 0.18 $\pm$ 0.01 0.01 $\pm$ 0.06 0.90 $\pm$ 0.06 0.80 $\pm$ 0.39 1.12 $\pm$ 0.03
gpt-4.1-mini 0.44 $\pm$ 0.02 0.54 $\pm$ 0.09 0.92 $\pm$ 0.02 0.10 $\pm$ 0.11 0.39 $\pm$ 0.00
gpt-4.1 0.47 $\pm$ 0.02 0.62 $\pm$ 0.06 0.97 $\pm$ 0.00 0.03 $\pm$ 0.13 0.53 $\pm$ 0.16
gpt-5-nano 0.31 $\pm$ 0.11 0.45 $\pm$ 0.16 0.89 $\pm$ 0.03 0.03 $\pm$ 0.13 0.06 $\pm$ 0.06
gpt-5-mini 0.47 $\pm$ 0.02 0.63 $\pm$ 0.07 0.93 $\pm$ 0.07 0.00 $\pm$ 0.00 0.03 $\pm$ 0.00
gpt-5 0.54 $\pm$ 0.03 0.88 $\pm$ 0.04 0.99 $\pm$ 0.03 0.00 $\pm$ 0.00 0.01 $\pm$ 0.04
@@
qwen3-14b 0.19 $\pm$ 0.01 0.49 $\pm$ 0.09 0.87 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
qwen3-32b 0.22 $\pm$ 0.01 0.55 $\pm$ 0.12 0.91 $\pm$ 0.08 0.05 $\pm$ 0.22 0.02 $\pm$ 0.04
qwen3-235b-a22b 0.41 $\pm$ 0.02 0.30 $\pm$ 0.00 0.73 $\pm$ 0.23 1.11 $\pm$ 0.55 0.03 $\pm$ 0.06
@@
gemini-2.0-flash-001 0.46 $\pm$ 0.02 0.67 $\pm$ 0.09 0.84 $\pm$ 0.00 0.00 $\pm$ 0.00 0.33 $\pm$ 0.07
gemini-2.5-flash-lite 0.28 $\pm$ 0.01 0.52 $\pm$ 0.04 0.91 $\pm$ 0.04 0.00 $\pm$ 0.00 0.57 $\pm$ 0.02
gemini-2.5-flash 0.46 $\pm$ 0.02 0.66 $\pm$ 0.16 0.95 $\pm$ 0.03 0.06 $\pm$ 0.13 0.59 $\pm$ 0.14
@@
claude-3.5-sonnet 0.55 $\pm$ 0.31 0.82 $\pm$ 0.06 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.39 $\pm$ 0.07
claude-sonnet-4.5 0.45 $\pm$ 0.02 0.80 $\pm$ 0.00 0.95 $\pm$ 0.04 0.00 $\pm$ 0.00 0.33 $\pm$ 0.00
@@
llama-3.1-70b-instruct 0.07 $\pm$ 0.04 0.63 $\pm$ 0.09 0.90 $\pm$ 0.05 0.77 $\pm$ 0.56 0.47 $\pm$ 0.07
llama-3.1-405b-instruct 0.18 $\pm$ 0.03 0.53 $\pm$ 0.16 0.86 $\pm$ 0.09 2.29 $\pm$ 0.25 0.57 $\pm$ 0.14
Average 0.37 0.57 0.91 0.28 0.35
Table 5: Psychometric contamination results for PVQ-40 with 95% confidence intervals from three experimental repetitions.
Semantic Memorization (Similarity $\uparrow$) Key Information Memorization (Success Rate $\uparrow$) Item–Dimension Mapping (F1 $\uparrow$) Option–Score Mapping (MAE $\downarrow$) (MAE $\downarrow$)
@@
gpt-4o-mini 0.21 $\pm$ 0.03 0.25 $\pm$ 0.00 0.93 $\pm$ 0.07 0.27 $\pm$ 0.14 0.60 $\pm$ 0.05
gpt-4o 0.24 $\pm$ 0.03 0.35 $\pm$ 0.12 1.00 $\pm$ 0.00 0.17 $\pm$ 0.38 0.73 $\pm$ 0.31
gpt-4.1-nano 0.19 $\pm$ 0.02 0.25 $\pm$ 0.10 0.94 $\pm$ 0.08 0.70 $\pm$ 0.38 1.46 $\pm$ 0.05
gpt-4.1-mini 0.26 $\pm$ 0.02 0.29 $\pm$ 0.10 0.87 $\pm$ 0.00 0.23 $\pm$ 0.52 0.79 $\pm$ 0.16
gpt-4.1 0.27 $\pm$ 0.04 0.40 $\pm$ 0.06 1.00 $\pm$ 0.00 0.33 $\pm$ 0.18 0.44 $\pm$ 0.13
gpt-5-nano 0.05 $\pm$ 0.05 0.32 $\pm$ 0.12 0.97 $\pm$ 0.03 0.00 $\pm$ 0.00 0.03 $\pm$ 0.05
gpt-5-mini 0.26 $\pm$ 0.03 0.51 $\pm$ 0.06 0.99 $\pm$ 0.05 0.00 $\pm$ 0.00 0.01 $\pm$ 0.03
gpt-5 0.70 $\pm$ 0.16 0.61 $\pm$ 0.06 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
@@
qwen3-14b 0.24 $\pm$ 1.39 0.22 $\pm$ 0.16 0.94 $\pm$ 0.05 0.07 $\pm$ 0.16 0.16 $\pm$ 0.20
qwen3-32b 0.19 $\pm$ 0.03 0.22 $\pm$ 0.06 0.99 $\pm$ 0.05 0.34 $\pm$ 0.39 0.12 $\pm$ 0.12
qwen3-235b-a22b 0.29 $\pm$ 0.05 0.17 $\pm$ 0.00 0.75 $\pm$ 0.17 0.86 $\pm$ 0.63 0.00 $\pm$ 0.00
@@
gemini-2.0-flash-001 0.22 $\pm$ 0.03 0.22 $\pm$ 0.06 0.64 $\pm$ 0.15 0.20 $\pm$ 0.00 0.50 $\pm$ 0.17
gemini-2.5-flash-lite 0.19 $\pm$ 0.03 0.24 $\pm$ 0.12 0.96 $\pm$ 0.00 0.40 $\pm$ 0.00 0.73 $\pm$ 0.16
gemini-2.5-flash 0.20 $\pm$ 0.03 0.28 $\pm$ 0.06 0.95 $\pm$ 0.05 0.10 $\pm$ 0.00 0.49 $\pm$ 0.04
@@
claude-3.5-sonnet 0.21 $\pm$ 0.14 0.51 $\pm$ 0.12 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.11 $\pm$ 0.04
claude-sonnet-4.5 0.11 $\pm$ 0.00 0.61 $\pm$ 0.06 0.95 $\pm$ 0.00 0.00 $\pm$ 0.00 0.66 $\pm$ 0.02
@@
llama-3.1-70b-instruct 0.05 $\pm$ 0.02 0.28 $\pm$ 0.16 0.86 $\pm$ 0.04 1.67 $\pm$ 0.04 0.67 $\pm$ 0.12
llama-3.1-405b-instruct 0.10 $\pm$ 0.08 0.18 $\pm$ 0.06 0.86 $\pm$ 0.08 2.23 $\pm$ 1.09 0.37 $\pm$ 0.09
Average 0.24 0.31 0.93 0.37 0.41
Table 6: Psychometric contamination results for SD-3 with 95% confidence intervals from three experimental repetitions.
Model Refusal Rate (%) Similarity Score (6-point scale)
OpenAI
gpt-4o-mini 0.0 $\pm$ 1.4 2.02 $\pm$ 0.28
gpt-4o 12.9 $\pm$ 5.7 1.77 $\pm$ 0.28
gpt-4.1-nano 9.1 $\pm$ 5.0 1.68 $\pm$ 0.25
gpt-4.1-mini 0.0 $\pm$ 1.4 1.72 $\pm$ 0.25
gpt-4.1 0.0 $\pm$ 1.4 1.92 $\pm$ 0.29
gpt-5-nano 62.3 $\pm$ 8.2 1.90 $\pm$ 0.51
gpt-5-mini 1.5 $\pm$ 2.5 3.07 $\pm$ 0.40
gpt-5 14.4 $\pm$ 6.0 5.74 $\pm$ 0.16
Qwen3
qwen3-14b 15.2 $\pm$ 6.1 1.59 $\pm$ 0.25
qwen3-32b 2.3 $\pm$ 2.8 1.52 $\pm$ 0.20
qwen3-235b-a22b 0.0 $\pm$ 1.8 1.69 $\pm$ 0.29
GLM
glm-4-32b 6.1 $\pm$ 4.2 1.66 $\pm$ 0.25
glm-4.5-air 2.3 $\pm$ 2.9 1.86 $\pm$ 0.26
glm-4.5 0.0 $\pm$ 1.4 3.32 $\pm$ 0.39
Gemini
gemini-2.0-flash-001 0.0 $\pm$ 1.4 1.64 $\pm$ 0.24
gemini-2.5-flash-lite 4.5 $\pm$ 3.7 1.67 $\pm$ 0.25
gemini-2.5-flash 0.0 $\pm$ 1.4 1.89 $\pm$ 0.30
Claude
claude-3.5-sonnet 98.5 $\pm$ 2.5 1.50 $\pm$ 6.35
claude-4.5-sonnet 71.2 $\pm$ 7.6 4.63 $\pm$ 0.65
Llama
llama-3.1-70b-instruct 87.9 $\pm$ 5.6 1.56 $\pm$ 0.82
llama-3.1-405b-instruct 97.7 $\pm$ 2.8 1.00 $\pm$ 0.00
Table 7: The refusal rate of each model and the semantic similarity score with 95% confidence intervals on BFI-44.
Model Refusal Rate (%) Similarity Score (6-point scale)
OpenAI
gpt-4o-mini 1.0 $\pm$ 2.7 1.19 $\pm$ 0.17
gpt-4o 99.0 $\pm$ 2.7 2.00 $\pm$ 0.00
gpt-4.1-nano 3.1 $\pm$ 3.9 1.29 $\pm$ 0.21
gpt-4.1-mini 0.0 $\pm$ 1.9 1.06 $\pm$ 0.09
gpt-4.1 0.0 $\pm$ 1.9 2.06 $\pm$ 0.34
gpt-5-nano 70.8 $\pm$ 9.0 1.50 $\pm$ 0.57
gpt-5-mini 2.1 $\pm$ 3.4 1.68 $\pm$ 0.32
gpt-5 41.1 $\pm$ 9.7 2.12 $\pm$ 0.49
Qwen3
qwen3-14b 26.0 $\pm$ 8.7 1.24 $\pm$ 0.23
qwen3-32b 5.2 $\pm$ 4.7 1.25 $\pm$ 0.20
qwen3-235b-a22b 0.0 $\pm$ 2.6 1.99 $\pm$ 0.36
GLM
glm-4-32b 0.0 $\pm$ 1.9 1.23 $\pm$ 0.17
glm-4.5-air 20.5 $\pm$ 8.9 1.16 $\pm$ 0.21
glm-4.5 6.2 $\pm$ 5.0 1.43 $\pm$ 0.25
Gemini
gemini-2.0-flash-001 0.0 $\pm$ 1.9 1.19 $\pm$ 0.19
gemini-2.5-flash-lite 40.6 $\pm$ 9.6 1.21 $\pm$ 0.21
gemini-2.5-flash 0.0 $\pm$ 1.9 1.43 $\pm$ 0.20
Claude
claude-3.5-sonnet 97.9 $\pm$ 3.4 1.00 $\pm$ 0.00
claude-4.5-sonnet 96.9 $\pm$ 3.9 6.00 $\pm$ 0.00
Llama
llama-3.1-70b-instruct 77.1 $\pm$ 8.3 1.36 $\pm$ 0.54
llama-3.1-405b-instruct 96.9 $\pm$ 3.9 1.00 $\pm$ 0.00
Table 8: The refusal rate of each model and the semantic similarity score with 95% confidence intervals on MFQ.
Model Refusal Rate (%) Similarity Score (6-point scale)
OpenAI
gpt-4o-mini 31.7 $\pm$ 5.8 1.55 $\pm$ 0.20
gpt-4o 42.5 $\pm$ 6.2 1.93 $\pm$ 0.27
gpt-4.1-nano 18.3 $\pm$ 4.9 1.56 $\pm$ 0.17
gpt-4.1-mini 0.0 $\pm$ 0.8 1.70 $\pm$ 0.18
gpt-4.1 0.0 $\pm$ 0.8 1.86 $\pm$ 0.20
gpt-5-nano 92.9 $\pm$ 3.3 1.71 $\pm$ 0.67
gpt-5-mini 9.6 $\pm$ 3.7 1.83 $\pm$ 0.19
gpt-5 16.0 $\pm$ 4.6 2.77 $\pm$ 0.29
Qwen3
qwen3-14b 15.1 $\pm$ 4.5 1.18 $\pm$ 0.10
qwen3-32b 19.2 $\pm$ 5.0 1.38 $\pm$ 0.15
qwen3-235b-a22b 0.0 $\pm$ 0.9 1.88 $\pm$ 0.22
GLM
glm-4-32b 3.8 $\pm$ 2.5 1.50 $\pm$ 0.15
glm-4.5-air 0.9 $\pm$ 1.5 1.61 $\pm$ 0.17
glm-4.5 0.4 $\pm$ 1.2 1.88 $\pm$ 0.21
Gemini
gemini-2.0-flash-001 0.0 $\pm$ 0.8 1.74 $\pm$ 0.18
gemini-2.5-flash-lite 0.8 $\pm$ 1.4 1.49 $\pm$ 0.15
gemini-2.5-flash 0.0 $\pm$ 0.8 1.78 $\pm$ 0.18
Claude
claude-3.5-sonnet 97.9 $\pm$ 1.9 2.00 $\pm$ 2.78
claude-4.5-sonnet 0.0 $\pm$ 0.8 1.98 $\pm$ 0.21
Llama
llama-3.1-70b-instruct 92.9 $\pm$ 3.3 1.00 $\pm$ 0.00
llama-3.1-405b-instruct 88.3 $\pm$ 4.1 1.21 $\pm$ 0.31
Table 9: The refusal rate of each model and the semantic similarity score with 95% confidence intervals on PVQ-40.
Model Refusal Rate Similarity Score (6-point scale)
OpenAI
gpt-4o-mini 0.0 $\pm$ 2.5 1.32 $\pm$ 0.23
gpt-4o 36.1 $\pm$ 10.8 1.54 $\pm$ 0.36
gpt-4.1-nano 12.5 $\pm$ 7.7 1.38 $\pm$ 0.25
gpt-4.1-mini 1.4 $\pm$ 3.6 1.90 $\pm$ 0.32
gpt-4.1 0.0 $\pm$ 2.5 2.22 $\pm$ 0.40
gpt-5-nano 91.7 $\pm$ 6.6 1.00 $\pm$ 0.00
gpt-5-mini 23.6 $\pm$ 9.7 2.60 $\pm$ 0.49
gpt-5 69.4 $\pm$ 10.4 4.18 $\pm$ 1.06
Qwen3
qwen3-14b 97.2 $\pm$ 4.4 3.00 $\pm$ 25.41
qwen3-32b 34.7 $\pm$ 10.7 1.53 $\pm$ 0.37
qwen3-235b-a22b 1.8 $\pm$ 4.6 2.47 $\pm$ 0.54
GLM
glm-4-32b 0.0 $\pm$ 2.5 1.53 $\pm$ 0.25
glm-4.5-air 19.4 $\pm$ 9.7 2.08 $\pm$ 0.47
glm-4.5 0.0 $\pm$ 2.5 2.17 $\pm$ 0.38
Gemini
gemini-2.0-flash-001 0.0 $\pm$ 2.5 1.88 $\pm$ 0.37
gemini-2.5-flash-lite 40.3 $\pm$ 11.0 1.30 $\pm$ 0.27
gemini-2.5-flash 1.4 $\pm$ 3.6 1.45 $\pm$ 0.26
Claude
claude-3.5-sonnet 93.1 $\pm$ 6.1 1.60 $\pm$ 1.67
claude-4.5-sonnet 98.6 $\pm$ 3.6 1.00 $\pm$ 0.00
Llama
llama-3.1-70b-instruct 69.4 $\pm$ 10.4 1.00 $\pm$ 0.00
llama-3.1-405b-instruct 95.8 $\pm$ 5.1 1.00 $\pm$ 0.00
Table 10: The refusal rate of each model and the semantic similarity score with 95% confidence intervals on SD-3.